prettydevil45:

input in standard form the equation of the given line. The line with m = -3/8 and passing through (5, -4)

prettydevil45:

@kaleighh533 @Joe2019

Joe2019:

idk

kaleighh533:

@darkknight

darkknight:

first use point slope form so y - y1 = m(x - x1) y - (-4) = -3/8(x- 5) not rewrite in slope-intercept form which is solving for y